OpenStudy (anonymous):

lim x->0 (1-x)^(2/x) i know that i have to use L'Hopitals Rule but im not sure how to move on from there. any help will be greatly appreciated

OpenStudy (turingtest):

do you know this property?\[\lim_{x\to a}f(x)=\lim_{x\to a}e^{\ln f(x)}\]?

OpenStudy (anonymous):

yes i did

OpenStudy (turingtest):

ok, so in a more convenient notation...\[\large\lim_{x\to 0}(1-x)^{\frac2 x}=\exp\left(\lim_{x\to 0}\ln(1-x)^{\frac2 x}\right)\]use log properties to simplify

OpenStudy (turingtest):

in case you haven't seen it, \[\exp(...)=e^{...}\]it's just a notation thing

OpenStudy (anonymous):

ok so now i would take the 2/x down so that i would get 2(1-x)/x?

OpenStudy (turingtest):

don't forget the ln

OpenStudy (anonymous):

so ln2(1-x)/x?

OpenStudy (turingtest):

the 2 goes in front of the ln

OpenStudy (turingtest):

\[\ln x^a=a\ln x\]

OpenStudy (anonymous):

2ln(1-x)/x?

OpenStudy (turingtest):

yes, that is what is in the exponent now use l'hospital

OpenStudy (anonymous):

i use l'hopitals rule just on the (1-x)/x?

OpenStudy (turingtest):

leave the 2 in there, but yes\[\large\lim_{x\to 0}(1-x)^{\frac2 x}=\exp\left(\lim_{x\to 0}\ln(1-x)^{\frac2 x}\right)=\exp\left(\lim_{x\to 0}{2\ln(1-x)\over x}\right)\]

OpenStudy (anonymous):

so the derivative will be -1/1 = -1 so i have e^2(-1)= 1/e^2?

OpenStudy (turingtest):

bingo :)

OpenStudy (anonymous):

thank you very much :D

OpenStudy (turingtest):

welcome!